NH2COONH4 when heated to 200 degree C gives a mixture of NH3 and CO2 vapours with a density of 16.0.


What is the degree of dissociation of NH2COONH4?


ans-1


please tell how?

Grade:12

3 Answers

vivek kumar
10 years ago

1

V. Harshit -
10 years ago

 

it is 1 because

NH2COONH4 when heated to 200 degree C gives a mixture of NH3 and CO2 vapours with a density of 16.0

Vikas TU
10 years ago

         NH2COONH4 --------> 2NH3 +  C02

t=0     a                                       0              0

t=t     a(1-k)                             2ak           ak

where k is the degree of diss. a is initial conc.

Now we know that vant hoff factor is given by total conc. at time ''t'' divided by total initial conc. i.e conc. of reactant initially. thus,

i = [a(1-k)  + 2ak + ak]/a

i = [a - ak +2ak + ak]/a

i = a(1+2k)/a

i = 1+ 2k

 

Now for strong electrolytes vant hoff factor ''i'' is equals to sum of no. of ions in the product.

therefore total no. of ions are = 2 +1 = 3

3 = 1 +2k

2k = 2

k =1

Hence the degree of dissociation for NH2COONH4 is equals to ''1''.

 

 

Plz Approve!